# Downscaling Neural Network for Coastal Simulations

### v2

@@ 5.4 Limitations and further discussion
One potential shortcoming of our proposed model is that it should be fine-tuned on other unknown regions for better performance. This is due to the domain-specific physical conditions at different places. For instance, the shoreline can affect fine-grained water movements. In our design, we define the landmass as zero, which serves as a hard condition for the downscaling neural network. One promising approach is to convert the coastal simulation into graphs and use graph neural networks to learn downscaling as graph interpolation. The advantage is that the water movement can pass along the vertices and edges and find nonlocal correlations for downscaling.
@@ 6 Conclusion
In this paper, we propose DNNCS , a spatiotemporal downscaling neural network for efficient coastal ocean simulation. The proposed spatiotemporal attention fully utilizes the neighborhood pixels across space and time via pixel shuffle to estimate the feature correlations. It can explicitly model the U-, V- and $\xi$-view changes over time as spatial feature correlations. The grid location is also embedded as the positional code for axis-aware attention estimation. Finally, we implement the novel feature split and reconstruction to project the features into the frequency domain for final enhancement. The initial spatiotemporal bilinear operator is used to encourage the network to learn the residues between LR and HR data. The proposed physics-informed losses also help to optimize the neural network with better quantitative results. In order to train the neural network, we propose a novel coastal simulation dataset on the Bahamas and Galveston. We use it for model training and evaluation. From extensive experiments and comparisons with other state-of-the-art methods, we can conclude that ours can achieve superior downscaling quality with fast computation.
Since the proposed methodology guarantees the mass conservation for arbitrarily long simulation times, it is particularly attractive for ocean, climate, and weather models for which high resolution simulations are desirable but computationally very expensive. In addition, it paves a novel direction for neural downscaling models for multigrid simulation, and we are interested in further exploration of ultra-resolution reconstruction and continuous downscaling.
@@ Acknowledgments
A. Rupp has been supported by the Academy of Finland’s grant number 350101 Mathematical models and numerical methods for water management in soils, grant number 354489 Uncertainty quantification for PDEs on hypergraphs, grant number 359633 Localized orthogonal decomposition for high-order, hybrid finite elements, Business Finland’s project number 539/31/2023 3D-Cure: 3D printing for personalized medicine and customized drug delivery, and the Finnish Flagship of advanced mathematics for sensing, imaging and modeling, decision number 358944. Markus Büttner was partially supported by the German Research Foundation (DFG) under grant AI 117/7-1 within the project 502500606 ’Performance-optimized co-design of ocean modeling software on FPGAs’.

### v3

@@ 5.6 Limitations and further discussion
A key limitation of the proposed framework is that optimal performance on previously unseen coastal regions generally requires fine-tuning with local data. This behavior arises from the strong domain-specific physical conditions inherent to different coastal environments, including shoreline geometry, bathymetry, and region-dependent hydrodynamic patterns, all of which influence fine-scale water movements. In practice, we identify two deployment modes. First, the pretrained model can be directly applied to a new region in a zero-shot manner when the land-sea mask is available. While this approach enables rapid, low-cost testing, its performance may vary depending on the degree of similarity between the source and target regions. Second, for operational use, fine-tuning the pretrained model with historical coarse–fine data pairs from the target region significantly improves accuracy by adapting the model to the local data distribution. Importantly, because the model is pretrained on physics-based simulations, the computational overhead of fine-tuning is modest, making this strategy feasible in real-world applications (see Section 5.5).
From a modeling perspective, our current design treats landmass as a hard constraint by assigning zero values over land, which effectively enforces physical boundaries but may limit the representation of complex coastal interactions near shorelines. A promising direction for future work is to reformulate coastal downscaling on irregular meshes as a graph-based learning problem. By representing coastal simulations as graphs and employing graph neural networks, water movement could propagate along vertices and edges, enabling the model to capture nonlocal dependencies and complex boundary interactions more naturally. Such an approach may further enhance domain transferability and robustness across diverse coastal regions.
@@ 6 Conclusion
In this paper, we propose DNNCS , a spatiotemporal downscaling neural network for efficient coastal ocean simulation. The proposed spatiotemporal attention fully utilizes the neighborhood pixels across space and time via pixel shuffle to estimate the feature correlations. It can explicitly model the U-, V- and $\xi$-view changes over time as spatial feature correlations. The grid location is also embedded as the positional code for axis-aware attention estimation. Finally, we implement the novel feature split and reconstruction to project the features into the frequency domain for final enhancement. The initial spatiotemporal bilinear operator is used to encourage the network to learn the residues between LR and HR data. The proposed physics-informed losses also help to optimize the neural network with better quantitative results. In order to train the neural network, we propose a novel coastal simulation dataset on the Bahamas and Galveston. We use it for model training and evaluation. From extensive experiments and comparisons with other state-of-the-art methods, we can conclude that ours can achieve superior downscaling quality with fast computation. In addition, we demonstrate the transferability of our methodology to other datasets and scenarios without extensive training, including flooding simulations not covered in our original training data.
Because the proposed methodology guarantees the mass conservation for arbitrarily long simulation times, it is particularly attractive for ocean, climate, and weather models for which high resolution simulations are desirable but computationally very expensive. In addition, it paves a novel direction for neural downscaling models for multigrid simulation, and we are interested in further exploration of ultra-resolution reconstruction and continuous downscaling.
Software and Data Availability
Software name: DNNCS — Developer: Zhi-Song Liu — First year available: 2025 — Hardware requirements: PC with Nvidia GPU — Software requirements: Pytorch, Pytorch Lightning, OpenCV, einops — Program language: Pytorch — Availability: https://github.com/Holmes-Alan/DNNCS — License: MIT — Test data and pretrained model: Same as the code repository
@@ Acknowledgments
A. Rupp has been supported by the Academy of Finland’s grants number 350101, 354489, 359633, 358944, Business Finland’s project number 539/31/2023, and Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) grant number 577175348. Markus Büttner was partially supported by DFG under grant AI 117/7-1 within the project 502500606 ‘Performance-optimized co-design of ocean modeling software on FPGAs’.
Bernhard Kainz acknowledges the HPC resources provided by NHR@FAU under the NHR projects b143dc and b180dc. NHR funding is provided by federal and Bavarian state authorities. NHR@FAU hardware is partially funded by the DFG - 440719683. Additional support was received by the ERC-project MIA-NORMAL 101083647, DFG 513220538, 512819079, and by the state of Bavaria (HTA).